OpenStudy (anonymous):

The following table shows the values of f(x) and g(x) for different values of x: x 1 2 3 4 5 f(x) 4 16 64 256 1024 g(x) -1 -1 -1 -1 -1 Which of the following best describes the function f(x) + g(x)? 4^x - 1x -1(4^x) 4^x - 1 4^x - 1

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

in case of g(x) its always -1 you now need to find f(x) as a formula hwo can we get 4 16 64 and 256 from x = 1 ,2 3 and 4?

OpenStudy (welshfella):

do you see the pattern.?

OpenStudy (anonymous):

Yes, you multiply by 4 every time

OpenStudy (anonymous):

Also, for the third option both the x and the -1 are exponents

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

yes you multiply by 4 to get the next term but what is the relation between each value of x and its corresponding value of f(x) how can we get from 1 to 4 2 to 16 3 to 64 using 1 formula

OpenStudy (welshfella):

look at the options for a clue 4 is involved in the formula which one of the first terms in the options looks likely?

OpenStudy (anonymous):

option d

OpenStudy (welshfella):

4^x is correct for f(x) as x = 1: 4^1 = 4 x = 2 4^2 = 16 x = 3 4^3 = 64 etc

OpenStudy (welshfella):

But c and d are the same . Looks like one is wrong

OpenStudy (anonymous):

Yes for c the hole thing, x-1, is the exponent

OpenStudy (welshfella):

4^x - 1 is the correct option. For g(x) its always = -1 whatever the value of x.

OpenStudy (welshfella):

OH OK well its d because you are adding: f(x) + g(x)

OpenStudy (welshfella):

OK?

OpenStudy (welshfella):

sorry but i gotta go right now.

OpenStudy (anonymous):

ok that's fine. thank you